An electric bulb marked 220 V, 100 W will get fused if it is made to consume 150 W or more. What voltage fluctuation will the bulb withstand?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. V = 220 v P = 100 w R = V^2/P = 220 * 220/100 = 484 Ω P = 150 w V = √PR = √150 * 22 *22 = 22√150 = 269.4 = 270 v